OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of fourteen shoes in his closet. The box contains 6 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be black? Round to the nearest percent if necessary. Choices: sixteen percent eight percent eighteen percent seven percent

OpenStudy (anonymous):

\[\frac{4}{14}\times \frac{3}{13}\]

OpenStudy (anonymous):

that is because it is "without replacement" with replacement it would be what luis rivera wrote